Charlene is evaluating a capital budgeting project that should last for 4 years. The project requires $950,000 of equipment and is eligible for 100% bonus depreciation. She is unsure whether immediately expensing the equipment or using straight-line depreciation is better for the analysis. Under straight-line depreciation, the cost of the equipment would be depreciated evenly over its 4-year life (ignore the half-year convention for the straight-line method). The company's WACC is 10%, and its tax rate is 30%.

  1. What would the depreciation expense be each year under each method? Enter your answers as positive values. Round your answers to the nearest dollar.
    Year Scenario 1
    (Straight-Line)
    Scenario 2
    (Bonus Depreciation)
    0 $   $  
    1 $   $  
    2 $   $  
    3 $   $  
    4 $   $  
  2. Which depreciation method would produce the higher NPV?
    -Select : Straight-Line or Bonus DepreciationItem
    How much higher would the NPV be under the preferred method? Do not round intermediate calculations. Round your answer to the nearest dollar.
    $  
Answer #1

Straight line depreciation = Original cost / Useful life

= 950000 / 4

= 237500

Hence depreciation from Year 1 to Year 4 = 237500

Under bonus method depreciation, depreciation at year 0 = 100% of initial investment = 950000

Part B) Under the 100% bonus method of depreciation, the NPV would be higher because the 100 % depreciation provides a tax shield today which will not be subject to any discounting. Whereas in straight line, each year 25% of depreciation tax shield will be discounted.

Depreciation tax saving under 100% bonus method = 950000 * tax rate = 950000 *30% = 285000

Under straight line, we need to calculate the present value of tax shield

Tax saving on depreciation annually = 237500 * 30% = 71250

Present value = 71250 * (1 - (1+r)^-n) /r

= 71250 * (1 - (1+10%)^-4)/10%

= 225852.91

Difference in NPV = 285000 - 225852.91 = 59147.09